OpenStudy (anonymous):

integral of (x^2-2x-1)/[(x-1)^2(x^2+1)]? I tried using rational functions with A/x-1, B/(x-1)^2, and (Cx+D)/(x^2+1)...

OpenStudy (anonymous):

for my systems I got: 0=A 0=-2A+C 1=2A+B-2C+D -2=-2A+C-2D -1=A+B+D

OpenStudy (anonymous):

they don't work out though